# Optimal Transmit Signal Design for Multi-Target MIMO Sensing Exploiting Prior Information

### v1

@@ Abstract
In this paper, we study the transmit signal optimization in a multiple-input multiple-output (MIMO) radar system for sensing the angle information of multiple targets via their reflected echo signals. We consider a challenging and practical scenario where the angles to be sensed are unknown and random, while their probability information is known a priori for exploitation. First, we establish an analytical framework to quantify the multi-target sensing performance exploiting prior distribution information, by deriving the posterior Cramér-Rao bound (PCRB) as a lower bound of the mean-squared error (MSE) matrix in sensing multiple unknown and random angles. Then, we formulate and study the transmit sample covariance matrix optimization problem to minimize the PCRB for the sum MSE in estimating all angles. By leveraging Lagrange duality theory, we analytically prove that the optimal transmit covariance matrix has a rank-one structure, despite the multiple angles to be sensed and the continuous feasible range of each angle. Moreover, we propose a sum-of-ratios iterative algorithm which can obtain the optimal solution to the PCRB-minimization problem with low complexity. Numerical results validate our results and the superiority of our proposed design over benchmark schemes.
@@ I Introduction
However, in practice, the location parameters to be sensed can be unknown and random, while the distribution of them can be known a priori based on e.g., target appearance pattern and statistical information. In this case, posterior Cramér-Rao bound (PCRB) or Bayesian Cramér-Rao bound (BCRB) can be adopted to characterize the lower bound of the MSE when the prior distribution information of the unknown and random parameters is exploited [8, 9, 10, 11, 12]. [8] studied the Bayesian sequential beamforming optimization in an ISAC system. [9] studied a MIMO radar system which aims to estimate the angle information of a point target. It was shown that the PCRB-minimizing transmit signal design will result in a novel probability-dependent power focusing effect. This work was extended to a MIMO ISAC system, where useful properties on the optimal transmit covariance matrix were derived [10]. [11, 12] considered a secure ISAC system where a multi-antenna base station (BS) communicates with one user and senses the location of a potential eavesdropping target simultaneously. [13] extended the exploitation of prior distribution information to the trajectory optimization of a sensing drone. Nevertheless, existing works along this line were focused on the sensing of a single target. With multiple targets, the optimal transmit signal design for optimizing the overall sensing performance of all targets still remains an open problem; moreover, how to exploit the prior joint distribution of all targets’ parameters in the design is also challenging. This thus motivates our study in this paper.
@@
We consider a MIMO radar system with multiple targets, whose unknown and random angles need to be sensed by exploiting their prior distribution information. To quantify the multi-target sensing performance, we derive the PCRB for the sensing MSE matrix as an explicit function of the transmit sample covariance matrix, which is general for any prior distribution. Based on this, we formulate the transmit sample covariance matrix optimization problem to minimize the PCRB for the sum MSE in sensing all targets’ angles, which is a convex optimization problem. By leveraging the Lagrange duality theory and Karush-Kuhn-Tucker (KKT) optimality conditions, we analytically prove that the optimal transmit covariance matrix has a rank-one structure, despite the existence of multiple targets and the continuous range of each target’s possible angle. Moreover, we propose a novel sum-of-ratios iterative algorithm to obtain the optimal solution with low complexity. It is shown via numerical results that our proposed design outperforms various benchmark designs.
@@ II System Model
Consider a MIMO radar system with $N_{t}\!\geq\!1$ transmit antennas and $N_{r}\!\!\geq\!\!1$ co-located receive antennas at the BS, which aims to sense the unknown and random angle information of $M\!\!\geq\!\!1$ point targets via the echo signals reflected by the targets. Let $\theta_{m}$ denote the angle of each $m$-th target with respect to the BS, as illustrated in Fig. 1. Let $p_{\boldsymbol{\Theta}}(\boldsymbol{\theta})$ denote the joint probability density function (PDF) of all angles in $\boldsymbol{\theta}=[\theta_{1},...,\theta_{M}]^{T}$, which is assumed to be known a priori via target appearance pattern or statistical information [9, 10, 12, 11].
@@ III-B Derivation of FIM from Prior Information, $\bm{F}_{\mathrm{P}}$
$\bm{F}_{\mathrm{P}}=\mathbb{E}_{\boldsymbol{\zeta}}\Big{[}\frac{\partial\ln\left(p_{\mathrm{Z}}\left(\boldsymbol{\zeta}\right)\right)}{\partial\boldsymbol{\zeta}}\big{(}\frac{\partial\ln\left(p_{\mathrm{Z}}\left(\boldsymbol{\zeta}\right)\right)}{\partial\boldsymbol{\zeta}}\big{)}^{H}\Big{]},\vspace{-0.15cm}$ (9)
@@ III-C Derivation of PCRB of MSE for Sensing $\bm{\theta}$
$\bm{F}^{-1}=\Big{[}\begin{matrix}\big{(}\bm{F}^{\mathrm{\bm{\theta\theta}}}_{\mathrm{O}}+\left[\bm{F}_{\mathrm{P}}\right]_{1:M,1:M}\big{)}^{-1}&\boldsymbol{C}\\ \boldsymbol{C}^{H}&\boldsymbol{D}\\ \end{matrix}\Big{]},\vspace{-0.15cm}$ (11)
@@ IV Problem Formulation
Note that Problem (P1) can be shown to be a convex optimization problem, since $\beta_{m}>0$ and $\bm{A}_{m}\succeq\bm{0}$. In the following, we study the structure of the optimal solution to unveil the fundamental transmit signal design principle for multi-target sensing exploiting prior information. Particularly, we aim to reveal the optimal rank of the transmit sample covariance matrix. Moreover, we will devise a low-complexity algorithm for obtaining the optimal solution to Problem (P1).
@@ V Optimal Solution Structure to Problem (P1)
In the following, we unveil the structure of the optimal $\boldsymbol{R}_{X}^{\star}$ by investigating (P2) via the Lagrange duality theory. Denote $\boldsymbol{Z}=[\boldsymbol{Z}_{1},...,\boldsymbol{Z}_{M}]$ with $\bm{Z}_{m}\succeq\bm{0},\forall m$, $\mu>0$, and $\boldsymbol{\varPsi}\succeq\bm{0}$ as the dual variables associated with the constraints in (17), (18), and (19), respectively. The Lagrangian of (P2) is given by
@@
where $\boldsymbol{P}_{m}\overset{\Delta}{=}[\beta_{m}\mathrm{tr}(\boldsymbol{A}_{m}\boldsymbol{R}_{X})+\delta_{m},1;1,t_{m}]$. By further defining $\boldsymbol{Z}_{m}=[z_{1,m},z_{2,m};z_{2,m}^{*},z_{3,m}]$, $\mathrm{tr}(\boldsymbol{Z}_{m}\boldsymbol{P}_{m})$ can be written as $\mathrm{tr}\left(\boldsymbol{Z}_{m}\boldsymbol{P}_{m}\right)=z_{1,m}\beta_{m}\mathrm{tr}\left(\boldsymbol{A}_{m}\boldsymbol{R}_{X}\right)+z_{1,m}\delta_{m}+z_{2,m}+z_{2,m}^{*}+z_{3,m}t_{m}$. Let $\boldsymbol{R}^{\star}_{X},\boldsymbol{t}^{\star},\boldsymbol{Z}^{\star},\mu^{\star},\boldsymbol{\varPsi}^{\star}$ denote the optimal primal and dual variables. The KKT optimality conditions consist of (17)-(19) and
@@
It follows from (21) that $z^{\star}_{3,m}=1$, which indicates $\boldsymbol{Z}^{\star}_{m}\neq\mathbf{0}$. Since $\boldsymbol{P}^{\star}_{m}\neq\mathbf{0}$, both $\boldsymbol{Z}^{\star}_{m}$ and $\boldsymbol{P}^{\star}_{m}$ should be singular such that (23) holds. Therefore, we have $z^{\star}_{1,m}-|z^{\star}_{2,m}|^{2}{z^{\star}_{3,m}}^{-1}=0$, and thus $z^{\star}_{1,m}=|z^{\star}_{2,m}|^{2}\geq 0$. Note that if $\bm{A}_{m}=\bm{0}$, any feasible $\bm{R}_{X}$ is optimal to (P1) and (P2), thus we only focus on the case of $\bm{A}_{m}\succ\bm{0}$, which can be shown to hold for various practical antenna array layouts such as uniform linear/planar array (ULA/UPA). With $\boldsymbol{A}_{m}\succ\mathbf{0}$ and $\boldsymbol{\varPsi}^{\star}\succeq\mathbf{0}$, we have $\mu^{\star}\neq 0$, thus $\mathrm{tr}\left(\boldsymbol{R}^{\star}_{X}\right)=P$. Based on (23), we further have
$\!\!\!\!t^{\star}_{m}=-(\mathrm{tr}(z^{\star}_{1,m}\beta_{m}\boldsymbol{A}_{m}\boldsymbol{R}^{\star}_{X})+z^{\star}_{1,m}\delta_{m}+z^{\star}_{2,m}+{z^{\star}_{2,m}}^{*}).\!\!\vspace{-0.1cm}$ (27)
@@
Based on (P2-eqv), we have the following proposition.
@@
For Problem (P1), an optimal transmit sample covariance matrix is $\bm{R}_{X}^{\star}=P\boldsymbol{q}_{1}\boldsymbol{q}_{1}^{H}$, where $\boldsymbol{q}_{1}$ is the eigenvector corresponding to the strongest eigenvalue of $\sum_{m=1}^{M}{\boldsymbol{H}^{\star}_{m}}$.
@@
Denote the eigenvalue decomposition (EVD) of $\sum_{m=1}^{M}{\boldsymbol{H}_{m}}^{\star}$ as $\boldsymbol{Q\varLambda Q}^{H}$, where $\bm{Q}=[\bm{q}_{1},...,\bm{q}_{N_{t}}]$ and $\boldsymbol{\varLambda}=\mathrm{diag}\{\lambda_{1},...,\lambda_{N_{t}}\}$ with $\lambda_{1}\geq...\geq\lambda_{N_{t}}$. Based on [9], there always exists a rank-one optimal solution to (P2-eqv) denoted by $\boldsymbol{R}_{X}^{\star}=P\boldsymbol{q}_{1}\boldsymbol{q}_{1}^{H}$. Proposition 1 thus follows based on the equivalence among (P1), (P2), and (P2-eqv). ∎
Proposition 1 indicates that although the angles of multiple targets need to be sensed, while each angle generally has a continuous region for its possible values, the optimal transmit sample covariance matrix for minimizing the PCRB for the sum MSE in estimating all these angles still has a rank-one structure. Moreover, the optimal transmit sample covariance matrix design is dependent on $\{\bm{A}_{m}\}_{m=1}^{M}$ which is determined by the joint PDF of all angles, $p_{\bm{\Theta}}(\bm{\theta})$, the array steering vectors at the BS transmit and receive antennas (i.e., the antenna layouts), as well as other system parameters including the transmit power budget, $P$, total number of samples used for sensing, $L$, and variances of the overall reflection coefficients for each target, $\{\sigma_{\alpha_{m}}^{2}\}_{m=1}^{M}$.
@@ VII Optimal Solution to Problem (P1) via Sum-of-Ratios Iterative Algorithm
where $\boldsymbol{\varphi}=[\varphi_{1},...,\varphi_{M}]^{T}\in\mathbb{R}^{M}$ is an auxiliary variable. Note that (P3) is a convex optimization problem which satisfies the Slater’s condition, thus strong duality holds. Denote $\boldsymbol{u}=[u_{1},...,u_{M}]^{T}\succeq\bm{0}$, $\mu>0$, and $\boldsymbol{\varPsi}\succeq\bm{0}$ as the dual variables associated with the constraints in (31), (32), and (33), respectively. The Lagrangian of (P3) is given by
@@
$-\sum_{m=1}^{M}{u_{m}^{\star}\varphi^{\star}_{m}\beta_{m}\boldsymbol{A}_{m}^{T}}+\mu\boldsymbol{I}_{N_{t}}-\boldsymbol{\varPsi}^{\star}$ $=\mathbf{0}$ (35)
@@
$\mathrm{tr}(\boldsymbol{\varPsi}^{\star}\bm{R}^{\star}_{X})$ $=\mathbf{0}$ (39)
@@
$u_{m}^{\star}=\left(\beta_{m}\mathrm{tr(}\boldsymbol{A}_{m}\boldsymbol{R}_{X}^{\star})+\delta_{m}\right)^{-1},\quad m=1,...,M.\vspace{-0.2cm}$ (41)
@@
$1-\varphi^{\star}_{m}(\beta_{m}\mathrm{tr}(\boldsymbol{A}_{m}\boldsymbol{R}^{\star}_{X})+\delta_{m})=0,\quad m=1,...,M.\vspace{-0.2cm}$ (42)
@@
s.t. $\mathrm{tr}(\boldsymbol{R}_{X})\leq P$ (47)
$\boldsymbol{R}_{X}\succeq\mathbf{0}.$ (48)
Denote $\sum_{m=1}^{M}{\!\!u_{m}\varphi_{m}\beta_{m}\boldsymbol{A}_{m}}\!\!=\!\tilde{\boldsymbol{Q}}\tilde{\boldsymbol{\varLambda}}\tilde{\boldsymbol{Q}}^{H}$, where $\tilde{\bm{Q}}\!\!=\!\![\tilde{\bm{q}}_{1},...,\tilde{\bm{q}}_{N_{t}}]$ and $\tilde{\boldsymbol{\varLambda}}=\mathrm{diag}\{\tilde{\lambda}_{1},...,\tilde{\lambda}_{N_{t}}\}$ with $\tilde{\lambda}_{1}\geq...\geq\tilde{\lambda}_{N_{t}}$. It can be shown that $\tilde{\boldsymbol{R}}_{X}^{\star}=P\tilde{\boldsymbol{q}}_{1}\tilde{\boldsymbol{q}}_{1}^{H}$ is an optimal solution to (P3-eqv’) which can be obtained via EVD. Then, with given $\bm{R}_{X}$, $(\boldsymbol{\varphi},\boldsymbol{u})$ can be updated via the modified Newton’s method [17]. By iteratively updating $\bm{R}_{X}$ and $(\boldsymbol{\varphi},\boldsymbol{u})$ until convergence, the optimal solution to (P3) and equivalently (P1) can be obtained [17]. The detailed steps of the proposed sum-of-ratios iterative algorithm are summarized in Algorithm 1. It is worth noting that the optimal solution to (P3-eqv’) in each iteration has a rank-one structure, which indicates that the converged optimal solution of $\bm{R}_{X}$ also has a rank-one structure. This is consistent with our analysis in Section V.

### v2

@@ Abstract
In this paper, we study the transmit signal optimization in a multiple-input multiple-output (MIMO) radar system for sensing the angle information of multiple targets via their reflected echo signals. We consider a challenging and practical scenario where the angles to be sensed are unknown and random, while their probability information is known a priori for exploitation. First, we establish an analytical framework to quantify the multi-target sensing performance exploiting prior distribution information, by deriving the posterior Cramér-Rao bound (PCRB) as a lower bound of the mean-squared error (MSE) matrix in sensing multiple unknown and random angles. Then, we formulate and study the transmit sample covariance matrix optimization problem to minimize the PCRB for the sum MSE in estimating all angles. Moreover, we propose a sum-of-ratios iterative algorithm which can obtain the optimal solution to the PCRB-minimization problem with low complexity. Numerical results validate our results and the superiority of our proposed design over benchmark schemes.
@@ I Introduction
However, in practice, the location parameters to be sensed can be unknown and random, while the distribution of them can be known a priori based on e.g., target appearance pattern and statistical information. In this case, posterior Cramér-Rao bound (PCRB) or Bayesian Cramér-Rao bound (BCRB) can be adopted to characterize the lower bound of the MSE when the prior distribution information of the unknown and random parameters is exploited [8, 9, 10, 11, 12, 13]. [8] studied the Bayesian sequential beamforming optimization in an ISAC system. [9] studied a MIMO radar system which aims to estimate the angle information of a point target. It was shown that the PCRB-minimizing transmit signal design will result in a novel probability-dependent power focusing effect. This work was extended to a MIMO ISAC system, where useful properties on the optimal transmit covariance matrix were derived [10]. [12, 13] considered a secure ISAC system where a multi-antenna base station (BS) communicates with one user and senses the location of a potential eavesdropping target simultaneously. [14] extended the exploitation of prior distribution information to the trajectory optimization of a sensing drone. Nevertheless, existing works along this line were focused on the sensing of a single target. With multiple targets, the optimal transmit signal design for optimizing the overall sensing performance of all targets still remains an open problem; moreover, how to exploit the prior joint distribution of all targets’ parameters in the design is also challenging. This thus motivates our study in this paper.
@@
We consider a MIMO radar system with multiple targets, whose unknown and random angles need to be sensed by exploiting their prior distribution information. To quantify the multi-target sensing performance, we derive the PCRB for the sensing MSE matrix as an explicit function of the transmit sample covariance matrix, which is general for any prior distribution. Based on this, we formulate the transmit sample covariance matrix optimization problem to minimize the PCRB for the sum MSE in sensing all targets’ angles, which is a convex optimization problem. Moreover, we propose a novel sum-of-ratios iterative algorithm to obtain the optimal solution with low complexity. It is shown via numerical results that our proposed design outperforms various benchmark designs.
@@ II System Model
Consider a MIMO radar system with $N_{t}\!\geq\!1$ transmit antennas and $N_{r}\!\!\geq\!\!1$ co-located receive antennas at the BS, which aims to sense the unknown and random angle information of $M\!\!\geq\!\!1$ point targets via the echo signals reflected by the targets. Let $\theta_{m}$ denote the angle of each $m$-th target with respect to the BS, as illustrated in Fig. 1. Let $p_{\boldsymbol{\Theta}}(\boldsymbol{\theta})$ denote the joint probability density function (PDF) of all angles in $\boldsymbol{\theta}=[\theta_{1},...,\theta_{M}]^{T}$, which is assumed to be known a priori via target appearance pattern or statistical information [9, 10, 11, 13, 12].
@@ III-B Derivation of FIM from Prior Information, $\bm{F}_{\mathrm{P}}$
$\bm{F}_{\mathrm{P}}=\mathbb{E}_{\boldsymbol{\zeta}}\Big{[}\frac{\partial\ln\left(p_{\mathrm{Z}}\left(\boldsymbol{\zeta}\right)\right)}{\partial\boldsymbol{\zeta}}\big{(}\frac{\partial\ln\left(p_{\mathrm{Z}}\left(\boldsymbol{\zeta}\right)\right)}{\partial\boldsymbol{\zeta}}\big{)}^{H}\Big{]},$ (9)
@@ III-C Derivation of PCRB of MSE for Sensing $\bm{\theta}$
$\bm{F}^{-1}=\Big{[}\begin{matrix}\big{(}\bm{F}^{\mathrm{\bm{\theta\theta}}}_{\mathrm{O}}+\left[\bm{F}_{\mathrm{P}}\right]_{1:M,1:M}\big{)}^{-1}&\boldsymbol{C}\\ \boldsymbol{C}^{H}&\boldsymbol{D}\\ \end{matrix}\Big{]},$ (11)
@@ IV Problem Formulation
Note that Problem (P1) can be shown to be a convex optimization problem, since $\beta_{m}>0$ and $\bm{A}_{m}\succeq\bm{0}$. In the following, we study the structure of the optimal solution to unveil the fundamental transmit signal design principle for multi-target sensing exploiting prior information. Moreover, we will devise a low-complexity algorithm for obtaining the optimal solution to Problem (P1).
@@ V Optimal Solution Structure to Problem (P1)
In the following, we unveil the structure of the optimal $\boldsymbol{R}_{X}^{\star}$ by investigating (P2) via the Lagrange duality theory. Denote $\boldsymbol{Z}=[\boldsymbol{Z}_{1},...,\boldsymbol{Z}_{M}]$ with $\bm{Z}_{m}\succeq\bm{0},\forall m$, $\mu\geq 0$, and $\boldsymbol{\varPsi}\succeq\bm{0}$ as the dual variables associated with the constraints in (17), (18), and (19), respectively. The Lagrangian of (P2) is given by
@@
where $\boldsymbol{P}_{m}\overset{\Delta}{=}[\beta_{m}\mathrm{tr}(\boldsymbol{A}_{m}\boldsymbol{R}_{X})+\delta_{m},1;1,t_{m}]$. By further defining $\boldsymbol{Z}_{m}=[z_{1,m},z_{2,m};z_{2,m}^{*},z_{3,m}]$, $\mathrm{tr}(\boldsymbol{Z}_{m}\boldsymbol{P}_{m})$ can be written as $\mathrm{tr}\left(\boldsymbol{Z}_{m}\boldsymbol{P}_{m}\right)=z_{1,m}\beta_{m}\mathrm{tr}\left(\boldsymbol{A}_{m}\boldsymbol{R}_{X}\right)+z_{1,m}\delta_{m}+z_{2,m}+z_{2,m}^{*}+z_{3,m}t_{m}$. Let $\boldsymbol{R}^{\star}_{X},\boldsymbol{t}^{\star},\boldsymbol{Z}^{\star},\mu^{\star},\boldsymbol{\varPsi}^{\star}$ denote the optimal primal and dual variables. The Karush-Kuhn-Tucker (KKT) optimality conditions consist of (17)-(19) and
@@
It follows from (21) that $z^{\star}_{3,m}=1$, which indicates $\boldsymbol{Z}^{\star}_{m}\neq\mathbf{0}$. Since $\boldsymbol{P}^{\star}_{m}\neq\mathbf{0}$, both $\boldsymbol{Z}^{\star}_{m}$ and $\boldsymbol{P}^{\star}_{m}$ should be singular such that (23) holds. Therefore, we have $z^{\star}_{1,m}-|z^{\star}_{2,m}|^{2}{z^{\star}_{3,m}}^{-1}=0$, and thus $z^{\star}_{1,m}=|z^{\star}_{2,m}|^{2}\geq 0$. With $\boldsymbol{A}_{m}\succeq\mathbf{0}$ and $\boldsymbol{\varPsi}^{\star}\succeq\mathbf{0}$, we have $\mu^{\star}\neq 0$, thus $\mathrm{tr}\left(\boldsymbol{R}^{\star}_{X}\right)=P$. Based on (23), we further have
$\!\!\!\!t^{\star}_{m}=-(\mathrm{tr}(z^{\star}_{1,m}\beta_{m}\boldsymbol{A}_{m}\boldsymbol{R}^{\star}_{X})+z^{\star}_{1,m}\delta_{m}+z^{\star}_{2,m}+{z^{\star^{*}}_{2,m}}).\!\!$ (27)
@@
Based on (P2-I), we have the following proposition.
@@
For Problem (P1), an optimal transmit sample covariance matrix is $\boldsymbol{R}_{X}^{\star}=\sum_{n=1}^{N}{P_{n}\boldsymbol{q}_{n}\boldsymbol{q}_{n}^{H}}$, where $\sum_{n=1}^{N}{P_{n}}=P$ and $\boldsymbol{q}_{n}$ is the eigenvector corresponding to the strongest eigenvalue of $\sum_{m=1}^{M}{\boldsymbol{H}^{\star}_{m}}$.
@@
Denote the eigenvalue decomposition (EVD) of $\sum_{m=1}^{M}{\boldsymbol{H}_{m}}^{\star}$ as $\boldsymbol{Q\varLambda Q}^{H}$, where $\bm{Q}=[\bm{q}_{1},...,\bm{q}_{N_{t}}]$ and $\boldsymbol{\varLambda}=\mathrm{diag}\{\lambda_{1},...,\lambda_{N_{t}}\}$ with $\lambda_{1}=...=\lambda_{N}>\lambda_{N+1}\geq...\geq\lambda_{N_{t}}$. The optimal solution to (P2-I) can be shown to have the structure of $\boldsymbol{R}_{X}^{\star}=\sum_{n=1}^{N}{P_{n}\boldsymbol{q}_{n}\boldsymbol{q}_{n}^{H}}$ with $\sum_{n=1}^{N}{P_{n}}=P$. Proposition 1 thus follows based on the equivalence between (P1) and (P2). ∎
Proposition 1 indicates that the optimal transmit sample covariance matrix design is dependent on $\{\bm{A}_{m}\}_{m=1}^{M}$ which is determined by the joint PDF of all angles, $p_{\bm{\Theta}}(\bm{\theta})$, the array steering vectors at the BS transmit and receive antennas (i.e., the antenna layouts), as well as other system parameters including the transmit power budget, $P$, total number of samples used for sensing, $L$, and variances of the overall reflection coefficients for each target, $\{\sigma_{\alpha_{m}}^{2}\}_{m=1}^{M}$.
@@ VII Optimal Solution to Problem (P1) via Sum-of-Ratios Iterative Algorithm
where $\boldsymbol{\varphi}=[\varphi_{1},...,\varphi_{M}]^{T}\in\mathbb{R}^{M}$ is an auxiliary variable. Note that (P3) is a convex optimization problem which satisfies the Slater’s condition, thus strong duality holds. Denote $\boldsymbol{u}=[u_{1},...,u_{M}]^{T}\succeq\bm{0}$, $\mu\geq 0$, and $\boldsymbol{\varPsi}\succeq\bm{0}$ as the dual variables associated with the constraints in (31), (32), and (33), respectively. The Lagrangian of (P3) is given by
@@
$-\sum_{m=1}^{M}{u_{m}^{\star}\varphi^{\star}_{m}\beta_{m}\boldsymbol{A}_{m}^{T}}+\mu^{\star}\boldsymbol{I}_{N_{t}}-\boldsymbol{\varPsi}^{\star}$ $=\mathbf{0}$ (35)
@@
$\mathrm{tr}(\boldsymbol{\varPsi}^{\star}\bm{R}^{\star}_{X})$ $=0$ (39)
@@
$u_{m}^{\star}=\left(\beta_{m}\mathrm{tr(}\boldsymbol{A}_{m}\boldsymbol{R}_{X}^{\star})+\delta_{m}\right)^{-1},\quad m=1,...,M.$ (41)
@@
$1-\varphi^{\star}_{m}(\beta_{m}\mathrm{tr}(\boldsymbol{A}_{m}\boldsymbol{R}^{\star}_{X})+\delta_{m})=0,\quad m=1,...,M.$ (42)
@@
s.t. $\mathrm{tr}(\boldsymbol{R}_{X})\leq P.$ (47)
Denote $\sum_{m=1}^{M}{\!\!u_{m}\varphi_{m}\beta_{m}\boldsymbol{A}_{m}}\!\!=\!\tilde{\boldsymbol{Q}}\tilde{\boldsymbol{\varLambda}}\tilde{\boldsymbol{Q}}^{H}$, where $\tilde{\bm{Q}}\!\!=\!\![\tilde{\bm{q}}_{1},...,\tilde{\bm{q}}_{N_{t}}]$ and $\tilde{\boldsymbol{\varLambda}}=\mathrm{diag}\{\tilde{\lambda}_{1},...,\tilde{\lambda}_{N_{t}}\}$ with $\tilde{\lambda}_{1}=...=\tilde{\lambda}_{N}>\tilde{\lambda}_{N+1}\geq...\geq\tilde{\lambda}_{N_{t}}$. It can be shown that $\tilde{\boldsymbol{R}}_{X}^{\star}=\sum_{n=1}^{N}{\tilde{P}_{n}\tilde{\boldsymbol{q}}_{n}\tilde{\boldsymbol{q}}_{n}^{H}}$ with $\sum_{n=1}^{N}{\tilde{P}_{n}}=P$ is an optimal solution to (P3-eqv’) which can be obtained via EVD. Then, with given $\bm{R}_{X}$, $(\boldsymbol{\varphi},\boldsymbol{u})$ can be updated via the modified Newton’s method [18]. By iteratively updating $\bm{R}_{X}$ and $(\boldsymbol{\varphi},\boldsymbol{u})$ until convergence, the optimal solution to (P3) and equivalently (P1) can be obtained [18]. The detailed steps of the proposed sum-of-ratios iterative algorithm are summarized in Algorithm 1.